OpenStudy (stormoshawty):

Simplify the equation. 3x^-2 / y

OpenStudy (stormoshawty):

@shamil98

OpenStudy (shamil98):

use the same rule. x^-n = 1/x^n what do you get for this?

OpenStudy (stormoshawty):

1/ x^2

OpenStudy (shamil98):

3/x^2* remember it's just 1/x^2 * 3 = 3/x^2 = 3x^-2 so 3/x^2 / y = 3/x^2 * 1/y = 3/x^(2)y